OpenStudy (brittany):

need help finishiing a problem x^2-10x+25=0 a=1 b=-10 c=25 10+-(square root)100-100/2

OpenStudy (anonymous):

\[(x-5)^2\] \[x=5\]

OpenStudy (anonymous):

a=1 b=-10 c=25 is correct (x-5)^2 = 0 x = 5

OpenStudy (anonymous):

Where'd you get stuck Brittany?

OpenStudy (brittany):

10±{-10^2-4(1)(25)}/2 {} stands for square root cos idont have a square root symbol

OpenStudy (brittany):

then i simplified and got 10±/2

OpenStudy (anonymous):

That's \[\frac{10 \pm 0}{2} = \frac{10}{2} = 5\]

OpenStudy (brittany):

and its only one answer? usually its like x = # or #

OpenStudy (anonymous):

Right, but in this case it's not because the part in the square root became a 0. Only if the square root part is non-zero do you get two solutions (one for the plus and one for the minus)

OpenStudy (brittany):

thankyou :)

OpenStudy (brittany):

polpak can you help me with http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e4970570b8b8d00ebe389b9